Point X is located at (2, −6) and point Z is located at (0, 5). Find the y value for the point Y that is located one over five the distance from point X to point Z.

Answer by stanbon(75887)

Point X is located at (2, −6) and point Z is located at (0, 5). Find the y value for the point Y that is located one over five the distance from point X to point Z.
--------
Let the point be (x,y)
x = (2-0)/5
y = (-6-5)/5
----
x = 2/5
y = -11/5 - -2 1/5
------------------
Cheers,
Stan H.
------------------